From a point on a straight road, two cars are driven in opposite directions, one at 58 miles per hour and the other at 62 miles per hour. In how many hours will they be 480 miles apart?
Found 2 solutions by nerdybill, checkley77:
Answer by nerdybill(7384)

From a point on a straight road, two cars are driven in opposite directions, one at 58 miles per hour and the other at 62 miles per hour. In how many hours will they be 480 miles apart?
.
You must apply the distance formula:
d = rt
where
d is distance
r is rate or speed
t is time
.
Let x = time driven
58x + 62x = 480
120x = 480
x = 4 hours

Answer by checkley77(12844)

58T+62T=480
120T=480
T=480/120
T=4 HOURS THEY WILL BE 480 MILES APART.